OpenStudy (anonymous):

solve (5/x+5)= (15/x+7)

OpenStudy (anonymous):

5/x-15/x=2 -10/x=2 -2x=10 x=-5

OpenStudy (anonymous):

you can use cross multiplication

OpenStudy (anonymous):

so multiply 5 by x+7 which is equal to 15 multiplied by x+5

OpenStudy (anonymous):

then that is 5(x+7)=15(x+5) which can be simplified to: 5x+35=15x+75

OpenStudy (anonymous):

after that i think u can can solve for x from there then make sure to check your answer by substituting the value of x into the original equation

OpenStudy (mayankdevnani):

\[\frac{5}{x+5}=\frac{15}{x+7}\]

OpenStudy (mayankdevnani):

\[5x+35=15x+75\]

OpenStudy (mayankdevnani):

\[-10x=40\]

OpenStudy (mayankdevnani):

\[x=-4\]

OpenStudy (anonymous):

ya you are right @mayankdevnani i was saying exactly the same thing except i explained my steps

OpenStudy (mayankdevnani):

yaa

OpenStudy (anonymous):

did you understand now @Kelly243563 ?

OpenStudy (mayankdevnani):

@pkjha3105 you are wrong

OpenStudy (anonymous):

ya i did not really understand you working @pkjha3105

OpenStudy (mayankdevnani):

@thivitaa

OpenStudy (anonymous):

Hahaha if I'm completely honest I'm so confused!

OpenStudy (anonymous):

um just read through the replies again

OpenStudy (anonymous):

Yeah thanks